# Scaling Offline Model-Based RL via Jointly-Optimized World-Action Model Pretraining

### v1

@@ Abstract
A significant aspiration of offline reinforcement learning (RL) is to develop a generalist agent with high capabilities from large and heterogeneous datasets. However, prior approaches that scale offline RL either rely heavily on expert trajectories or struggle to generalize to diverse unseen tasks. Inspired by the excellent generalization of world model in conditional video generation, we explore the potential of image observation-based world model for scaling offline RL and enhancing generalization on novel tasks. In this paper, we introduce JOWA: Jointly-Optimized World-Action model, an offline model-based RL agent pretrained on multiple Atari games to learn general-purpose representation and decision-making ability. Our method jointly optimizes a world-action model through shared transformer backbone, which stabilize temporal difference learning with large models during pretraining. Moreover, we propose an provably efficient and parallelizable planning algorithm to compensate for the Q-value estimation error and thus search out better policies. Experimental results indicate that our largest agent, with 150 million parameters, achieves 78.9% human-level performance on pretrained games using only 10% subsampled offline data, outperforming existing state-of-the-art large-scale offline RL baselines by 31.6% on averange. Furthermore, JOWA scales favorably with model capacity and can sample-efficiently transfer to novel games using only 5k offline fine-tuning data corresponding to about 4 trajectories per game, which demonstrates superior generalization of JOWA. We will release codes at https://github.com/CJReinforce/JOWA.
@@ 1 Introduction
In recent years, building large-scale generalist models capable of solving multiple tasks has become a dominant research focus in natural language processing (NLP) and multi-modality. Notable examples include large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Team et al., 2023) and large vision-language models (Zhu et al., 2023; Liu et al., 2024; Bai et al., 2023), which deliver outstanding performance across a wide range of tasks and adapt quickly to new ones through few-shot or in-context learning. Their success is largely driven by the scaling law (Kaplan et al., 2020), which posits that increasing model size and data leads to improved performance. However, similar scaling trends have not been extensively observed in reinforcement learning (RL).
@@
Unlike in vision and language domains, RL has traditionally favored smaller models tailored to single tasks or multiple tasks within the same environment. Concerningly, previous studies have shown that scaling model capacity can lead to instabilities or performance degradation (Kumar et al., 2020a; Ota et al., 2021; Sokar et al., 2023), explaining the continued dominance of shallow CNN networks in vision-based RL tasks (Mnih, 2013). While some efforts have been made to scale offline RL across multiple tasks (Lee et al., 2022; Reed et al., 2022; Xu et al., 2022; Wu et al., 2024), they predominantly rely on supervised learning (SL) approaches, such as conditional behavior cloning, rather than temporal difference (TD) learning, and heavily rely on large amounts of expert trajectories. Kumar et al. (2023) scaled offline Q-learning using ResNet-based representation network with separate Q-heads for each game, but this approach only learns generalizable representations and still requires substantial data and gradient steps to adapt to unseen games due to reset of Q-heads during fine-tuning. Therefore, scaling TD-based offline RL for simultaneous general-purpose representation and decision-making remains a critical challenge. While Hansen et al. (2024) attempted to address this challenge for continuous control tasks with low-dimensional proprioceptive states using model-based RL, it lacks generalization due to heterogeneous proprioceptors across task domains.
@@
Meanwhile, world model has decoupled from model-based RL and evolved into a distinct research area within computer vision and multi-modality, primarily focusing on conditional video generation (Hong et al., 2022; Blattmann et al., 2023; Yu et al., 2023; Yang et al., 2024). Notably, SORA (Brooks et al., 2024) has demonstrated superior generalization performance as a world simulator through large-scale training of generative models on time-series image data. This motivates our investigation into a compelling question: “Can image observation-based world model scale offline RL across multiple tasks while enhancing generalization to diverse unseen tasks?
@@
To address this question, we introduce JOWA: Jointly-Optimized World-Action model, an offline model-based RL agent pretrained across multiple visual games. Crucially, JOWA unlocks scaling trends and achieves sample-efficient adaption to novel games. By utilizing a shared transformer backbone for both the world model and Q-value criticism, JOWA learns both generalizable representations and decision-making skills. This architecture allows the transformer to absorb gradients back-propagated from both world modeling and TD losses, enabling joint optimization. The world modeling loss acts as a regularizer, stabilizing TD-learning for large models. Additionally, we propose a provably efficient and parallelizable planning algorithm to compensate for the Q-value estimation error, allowing for consistent identification of the optimal policy at inference time and sample-efficient transfer to novel games.
@@
To evaluate the performance of JOWA, we train a single model to play 15 Atari games, similar to Lee et al. (2022); Kumar et al. (2023) but using a reduced yet sufficient dataset of 10M transitions per game—termed as the low-data regime to highlight the data efficiency. This setup presents a significant challenge due to the diverse nature of games, each with unique dynamics, rewards, visuals, and agent embodiments. To further test JOWA’s generalization ability, we perform offline fine-tuning on 5 unseen games, using minimal fine-tuning data.
@@
Our contributions are threefold: First, we introduce JOWA, an offline model-based RL method capable of training a single high-performing generalist agent across multiple Atari games. JOWA attains 78.9% human-level performance on pretrained games using only 10% of the original dataset (Agarwal et al., 2020), outperforming existing state-of-the-art large-scale offline RL baselines by 31.6% on averange. Second, we demonstrate that JOWA unlocks scaling trends, with performance improving as model capacity increases. Third, JOWA enables sample-efficient transfer to diverse unseen games with 64.7% DQN-normalized score using only 5k transitions per game. Our ablation studies highlight the significance of two key design features of JOWA: joint optimization and planning, along with other training choices. Due to the extensive training time, we will release all training, evaluation, and fine-tuning codes, as well as model weights, to support future research.
@@ 2 Related work
Offline RL algorithms learn a policy entirely from the static offline dataset without online interactions. Model-free offline RL incorporates conservatism to mitigate extrapolation error (Jin et al., 2021) primarily through policy constraints (Fujimoto et al., 2019; Kumar et al., 2019; Peng et al., 2019; Nair et al., 2020; Kostrikov et al., 2021; Fujimoto & Gu, 2021) and value regularization (Kumar et al., 2020b). In contrast, model-based offline RL approximates the environment using learned world models and performs conservative policy optimization (Lu et al., 2021; Yu et al., 2020b; 2021; Schrittwieser et al., 2021). While these works focus on single-task settings, our work explores scaling offline model-based RL across diverse, challenging multi-task Atari games (Lee et al., 2022; Kumar et al., 2023; Wu et al., 2024) aiming for sample-efficient transfer to novel games. We employ CQL (Kumar et al., 2020b) to estimate conservative Q-values, leveraging its simplicity and effectiveness on vision-based offline RL tasks.
@@
Multi-task reinforcement learning (MTRL) aims to learn a shared policy for diverse tasks, with various approaches proposed in the literature (Teh et al., 2017; Song et al., 2019; Xu et al., 2020; Yang et al., 2020; Sodhani et al., 2021). A common approach is to formulate the multi-task model as task-condition (Yu et al., 2020a), such as language-conditioned tasks (Lynch & Sermanet, 2020; Ahn et al., 2022; Jang et al., 2022) and goal-conditional RL (Plappert et al., 2018). In multi-task offline RL, conditional sequence modeling approaches based on decision transformer (Chen et al., 2021) or diffusion model (Janner et al., 2022) typically rely on large amounts of expert trajectories (Reed et al., 2022; Xu et al., 2022; Lee et al., 2022; Wu et al., 2024; Hu et al., 2024; He et al., 2024). Beyond conditional behavior cloning (BC) methods, Scaled-QL (Kumar et al., 2023) scales offline Q-learning using a shared feature network across tasks with separate Q-value heads for each task. Our work advances offline TD learning to multi-task settings without task-specific Q-value heads through a jointly-optimized world-action model. Table 1 compares experimental environments and open-source status of multi-task offline RL algorithms. Following Lee et al. (2022); Kumar et al. (2023); Wu et al. (2024), we focus on the multi-game regime, which presents greater challenges due to high-dimensional observations and diverse, stochastic environment dynamics. Due to the prohibitive long training time, we advocate for open-sourcing codes and checkpoints to facilitate progress in this field.
Table 1: Comparison of methods in multi-task offline RL. $\clubsuit$ and $\spadesuit$ represent two training paradigms of agents: conditional BC and TD-learning, respectively.
Method Experimental environment Open source code
@@ 3.2 Value regularization based Offline rl (CQL)
COMBO (Yu et al., 2021), a model-based variant of CQL, uses the following loss function:
$\mathcal{L}_{\text{COMBO}}=\alpha(\mathbb{E}_{s,a\sim\rho_{\hat{\mathcal{M}}}(s,a)}$ $\left[Q_{\theta}(s,a)\right]-\mathbb{E}_{s,a\sim\mathcal{D}}\left[Q_{\theta}(s,a)\right])$
$+$ $\frac{1}{2}\mathbb{E}_{s,a,s^{\prime}\sim d_{f}}\left[\left(\mathcal{T}^{*}Q_{\theta}(s,a)-Q_{\theta}(s,a)\right)^{2}\right]$ (4)
where $\hat{\mathcal{M}}$ is the imagined Markovian Deicision Process induced by the world model. $\rho_{\hat{\mathcal{M}}}(s,a)$ is the occupancy measure of current policy within $\hat{\mathcal{M}}$. $d_{f}$ is an $f$-interpolation between the offline dataset and synthetic rollouts from the world model: $d_{f}=fd_{\mathcal{D}}+(1-f)d_{\hat{\mathcal{M}}}$, where $f\in[0,1]$.
@@ 3.3 Problem setup
We consider a multi-task offline RL problem: given a static dataset of transitions $\mathcal{D}=\{(s_{t},a_{t},r_{t},d_{t},s_{t+1})_{i}\}$ collected from various environments with arbitrary behaviour polices, our goal is to learn a single policy that maximize the expected return $\mathcal{R}_{t}=\sum_{k\geq t}\gamma^{k-t}r_{k}$ on all considered environments and can be efficiently fine-tuned to new tasks, where $\gamma$ is the discount factor.
Considering the computational demands and resource limitations, we use a subset of 15 games from Atari 2600 for training. The whole training process took approximately 12 days on A100 GPUs. Our offline dataset is derived from the the DQN-Replay dataset (Agarwal et al., 2020), which consists of $84\times 84$ grayscale images as observations and a full action space of 18 discrete actions.
@@ 4 Jointly-Optimized World-Action Model
In this section, we first detail the architecture of JOWA and the loss functions for joint optimization in section 4.1. Next, we introduce the provably efficient and parallelizable planning algorithm employed to compensate for the Q-value estimation error in section 4.2. Finally, we present the overall training and sample-efficient fine-tuning pipelines in section 4.3.
@@ 4.1.2 Training of World-part Module
To unify the type of loss functions for balanced training (Vandenhende et al., 2021), we convert scalar rewards to ternary quantities $\{-1,0,1\}$ using the sign function. This allows all three predictors to be optimized as classification problems using cross-entropy loss. Given $L$-timesteps segments sampled from the offline dataset, the loss function for the world-part module is formulated as:
@@ 4.1.3 Training of Action-part Module
We use CQL (Kumar et al., 2020b) for offline TD-learning. To ensure training stability and enhance scaling performance (Farebrother et al., 2024), we employ distributional TD-error (Bellemare et al., 2017) instead of the mean-square TD-error, maintaining consistency with the $\mathcal{L}_{\text{world}}$ loss type.
The action-part module computes the return distribution for all actions given an observation and historical information. For an observation-action pair $(o_{t},a_{t})$, the return distribution $\mathcal{Z}$ is formulated as: $\mathcal{Z}(o_{t},a_{t})=h_{Q}(f(z_{\leq t},a_{<t},u))[a_{t}]$. The value function $Q(o_{t},a_{t})$ is the mean of $\mathcal{Z}(o_{t},a_{t})$. Then the loss function for the action-part module is formulated as:
@@
where $\mathcal{T}^{*}$ is the distributional Bellman optimality operator defined in Equation (1), and $\mathcal{Z}^{-}$ is the target distribution computed through a target Q-values head $h_{Q^{-}}$.
@@ 4.2 Parallelizable planning at inference time
Given an estimated optimal Q-value $\hat{Q}^{*}$, a learned world model with dynamic predictor $\hat{P}$ and reward predictor $\hat{r}$, our objective is to find the optimal action $a^{*}_{0}$ maximizing the ground-truth optimal Q-value $Q^{*}$, starting from initial state $s_{0}$. To do so, we rewrite the Bellman optimality equation as:
@@
where $\pi_{Q^{*}}$ is the policy induced by the optimal Q-function. The proof is provided in the Appendix A. For Equation (11), the optimal policy is the greedy policy based on $Q^{*}$.
@@
We implement this search using beam search, an efficient decoding algorithm common in NLP. At each timestep, we retain only $K$ states with the top-$K$ total income values. The horizon $H$ and beam width $K$ are hyper-parameters, with $K=1$ or $H=0$ degenerating to a $\hat{Q}^{*}$-based greedy policy. The algorithm calls the world model $K^{2}(H-1)+K$ times but only takes $H$ times as long as the forward propagation of the world model due to parallelizability between $K$ beams.
4.3 Training and Fine-tuning pipelines
@@
For sample-efficient transfer to unseen games, we adopt a similar two-stage offline fine-tuning pipeline, but unfreeze all components of JOWA in both two stages. In stage 1, we fine-tune JOWA using real data for 3k steps. In stage 2, we enable planning to synthesize high-quality data, using 3/4 batch of real and 1/4 batch of synthetic data with COMBO loss for 7k steps. Detailed fine-tuning protocol is shown in Appendix E.2.
@@ 5.1 Experimental Setup
We use the Atari trajectory dataset from Agarwal et al. (2020), which contains 50M transitions from each of 5 separate training runs. Due to the prohibitive long training time on full games, we select a subset of 20 games, maintaining the difficulty distribution of full games defined by Gulcehre et al. (2020). These 20 games are introduced in Appendix D. 15 of those games are used for training, and 5 games are held out for OOD generalization experiments. Following Lee et al. (2022); Kumar et al. (2023); Wu et al. (2024), we use data from 2 out of 5 training runs. To investigate performance in low-data regime, we uniformly draw 10% of transitions at random, as per Agarwal et al. (2020); Kumar et al. (2020a), resulting in 10M transitions per game.
@@
We implement our world-action model based on GPT-2 (Brown et al., 2020). We train a total of three variants of JOWA: JOWA-150M (150M parameters), JOWA-70M, and JOWA-40M. We set the number of visual tokens $K$ to 36 and sequence length $L$ to 8 game frames, resulting in sequences of 296 tokens. We pretrain all JOWA models on A100 GPUs for 1.75M steps using the AdamW optimizer (Loshchilov & Hutter, 2018) with learning rates of $1\cdot 10^{-4}$ and $5\cdot 10^{-5}$ for two stages respectively, and a batch size of 512 for both stages. For fine-tuning on novel games, we train for 10k gradient steps with a batch size of 32 and 5k transitions. We compare JOWA with the following baselines: (i) Multi-Task BC (MTBC) (ii) Multi-Game DT (MGDT) (Lee et al., 2022) (iii) Elastic DT (EDT) Wu et al. (2024) (iv) Scaled-QL (SQL) Kumar et al. (2023). For fair comparison, all methods use the same batch size of 512 and 1.75M gradient steps. The implementation details of all baselines are provided in the Appendix E.1.
@@
During evaluation, we set the inverse temperature $\kappa$ to 10 for EDT and MGDT for expert action inference. We enable planning for JOWA, setting the planning horizon $H$ to 2 for all games and adjust the beam width $K$ base on the size of valid action space of each game. See Appendix E.3 for the detailed evaluation protocol. We measure performance using human normalized scores (HNS) (Mnih et al., 2015), i.e. $(\textrm{score}-\textrm{score}_{\textrm{random}})/(\textrm{score}_{\textrm{human}}-\textrm{score}_{\textrm{random}})$. To create an aggregate comparison metric across all games, we use inter-quartile mean (IQM) of human-normalized scores, following evaluation best practices proposed in (Agarwal et al., 2021).

### v3

@@ Abstract
A significant aspiration of offline reinforcement learning (RL) is to develop a generalist agent with high capabilities from large and heterogeneous datasets. However, prior approaches that scale offline RL either rely heavily on expert trajectories or struggle to generalize to diverse unseen tasks. Inspired by the excellent generalization of world model in conditional video generation, we explore the potential of image observation-based world model for scaling offline RL and enhancing generalization on novel tasks. In this paper, we introduce JOWA: Jointly-Optimized World-Action model, an offline model-based RL agent pretrained on multiple Atari games with 6 billion tokens data to learn general-purpose representation and decision-making ability. Our method jointly optimizes a world-action model through a shared transformer backbone, which stabilize temporal difference learning with large models during pretraining. Moreover, we propose a provably efficient and parallelizable planning algorithm to compensate for the Q-value estimation error and thus search out better policies. Experimental results indicate that our largest agent, with 150 million parameters, achieves 78.9% human-level performance on pretrained games using only 10% subsampled offline data, outperforming existing state-of-the-art large-scale offline RL baselines by 71.4% on averange. Furthermore, JOWA scales favorably with model capacity and can sample-efficiently transfer to novel games using only 5k offline fine-tuning data (approximately 4 trajectories) per game, demonstrating superior generalization. The code and checkpoints will be released at https://github.com/CJReinforce/JOWA.
@@ 1 Introduction
In recent years, building large-scale generalist models capable of solving multiple tasks has become a dominant research focus in natural language processing (NLP) and multi-modality (Yue et al., 2024). Their success is largely driven by the scaling law (Kaplan et al., 2020), which posits that increasing model size and data leads to improved performance. However, similar scaling trends have not been extensively observed in reinforcement learning (RL).
@@
Unlike in vision and language domains, RL has traditionally favored smaller models tailored to single tasks or multiple tasks within the same environment. Concerningly, previous studies have shown that scaling model capacity can lead to instabilities or performance degradation (Kumar et al., 2020a; Ota et al., 2021; Sokar et al., 2023). While some efforts have been made to scale offline RL across multiple tasks (Lee et al., 2022; Xu et al., 2022), they predominantly rely on supervised learning (SL) approaches, such as conditional behavior cloning, rather than temporal difference (TD) learning, and heavily rely on large amounts of expert trajectories. Kumar et al. (2023) scaled offline Q-learning using ResNet-based representation network with separate Q-heads for each game, but this approach only learns generalizable representations and still requires substantial data and gradient steps to adapt to unseen games due to reset of Q-heads during fine-tuning. Therefore, scaling TD-based offline RL for simultaneous general-purpose representation and decision-making remains a critical challenge. While Hansen et al. (2024) attempted to address this challenge for continuous control tasks with low-dimensional proprioceptive states using model-based RL, it lacks generalization due to heterogeneous proprioceptors across task domains.
@@
Meanwhile, world model has decoupled from model-based RL and evolved into a distinct research area within computer vision and multi-modality, primarily focusing on conditional video generation (Hong et al., 2022; Blattmann et al., 2023; Yu et al., 2023; Yang et al., 2024; Bruce et al., 2024). Notably, SORA (Brooks et al., 2024) has demonstrated superior generalization performance as a world simulator through large-scale training of generative models on time-series image data. This motivates our investigation into a compelling question: “Can image observation-based world model scale offline RL across multiple tasks while enhancing generalization to diverse unseen tasks?
@@
To address this question, we introduce JOWA: Jointly-Optimized World-Action model, an offline model-based RL agent pretrained across multiple visual games with approximately 6 billion tokens data. Crucially, JOWA unlocks scaling trends and achieves sample-efficient adaption to novel games. By utilizing a shared transformer backbone for both the world model and Q-value criticism, JOWA learns both generalizable representations and decision-making skills. This architecture allows the transformer to absorb gradients back-propagated from both world modeling and TD losses, enabling joint optimization. The world modeling loss acts as a regularizer, stabilizing TD-learning for large models. Additionally, we propose a provably efficient and parallelizable planning algorithm to compensate for the Q-value estimation error, allowing for consistent identification of the optimal policy at inference time and sample-efficient transfer to novel games.
@@
To evaluate the performance of JOWA, we train a single model to play 15 Atari games, similar to Lee et al. (2022) but using a reduced yet sufficient dataset of 10M transitions per game—termed as the low-data regime to highlight the data efficiency. This setup presents a significant challenge due to the unique dynamics, visuals, and agent embodiments of games. To further test JOWA’s generalization, we perform offline fine-tuning on 5 unseen games, using minimal fine-tuning data.
@@
Our contributions are threefold: First, we introduce JOWA, an offline model-based RL method capable of training a single high-performing generalist agent across multiple Atari games. JOWA attains 78.9% human-level performance on pretrained games using only 10% of the original dataset (Agarwal et al., 2020), outperforming existing state-of-the-art large-scale offline RL baselines by 71.4% on averange. Second, we demonstrate that JOWA unlocks scaling trends, with performance improving as model capacity increases. Third, JOWA enables sample-efficient transfer to diverse unseen games with 64.7% DQN-normalized score using only 5k transitions per game, surpassing baselines by 69.9% on average. Our ablation studies highlight the significance of two key design features of JOWA: joint optimization and planning, along with other training choices. We release all the training, evaluation, and fine-tuning code, as well as model weights, to support future research.
@@ 2 Related work
Offline RL algorithms learn a policy entirely from the static offline dataset without online interactions. Model-free offline RL incorporates conservatism to mitigate extrapolation error (Jin et al., 2021) primarily through policy constraints (Fujimoto et al., 2019; Kumar et al., 2019; Fujimoto & Gu, 2021; Cheng et al., 2024; Liu et al., 2024; 2025) and value regularization (Kumar et al., 2020b). Model-based offline RL approximates the environment using world models and performs conservative policy optimization (Yu et al., 2020; 2021). While these works focus on single-task settings, our work explores scaling offline model-based RL across diverse, challenging multi-task Atari games (Lee et al., 2022; Kumar et al., 2023; Wu et al., 2024) aiming for sample-efficient transfer to novel games.
@@
Multi-task reinforcement learning (MTRL) aims to learn a shared policy for diverse tasks. A common approach is to formulate the multi-task model as task-condition, such as language-conditioned tasks (Ahn et al., 2022; Jang et al., 2022) and goal-conditional RL (Plappert et al., 2018). In multi-task offline RL, conditional sequence modeling approaches based on decision transformer (Chen et al., 2021) or diffusion model (Janner et al., 2022) typically rely on large amounts of expert trajectories (Xu et al., 2022; Lee et al., 2022; Wu et al., 2024; Hu et al., 2024; He et al., 2024). Beyond that, FICC (Ye et al., 2022) pretrains multi-task representation and dynamic models with action-free videos and then fine-tunes a model-based agent on each task for fast adaptation. Scaled-QL (Kumar et al., 2023) scales offline Q-learning using a shared feature network across tasks with separate Q-value heads for each task. Our work advances offline TD learning to multi-task settings without task-specific Q-value heads through a jointly-optimized world-action model. Table 1 compares experimental environments and open-source status of multi-task offline RL algorithms. Following Lee et al. (2022); Kumar et al. (2023); Wu et al. (2024), we focus on the multi-game regime, which presents greater challenges due to high-dimensional observations and diverse, stochastic environment dynamics.
Table 1: Comparison of methods in multi-task offline RL. $\clubsuit$ and $\spadesuit$ represent two training paradigms of agents, conditional BC and TD-learning, respectively.
Method Experimental environment Open source
@@ 3.3 Problem setup
We consider a multi-task offline RL problem: given a static dataset of transitions $\mathcal{D}=\{(s_{t},a_{t},r_{t},d_{t},s_{t+1})_{i}\}$ collected from various environments with arbitrary behaviour polices, our goal is to learn a single policy that maximizes the expected return $\mathcal{R}_{t}=\sum_{k\geq t}\gamma^{k-t}r_{k}$ on all considered environments and can be efficiently fine-tuned to new tasks, where $\gamma$ is the discount factor.
Considering the computational budget, we use 15 Atari games for pretraining and 5 games for out-of-distribution (OOD) experiments. The whole training process took around 12 days on A100 GPUs. Our offline dataset is derived from the the DQN-Replay dataset (Agarwal et al., 2020), which consists of $84\times 84$ grayscale images as observations and a full action space with 18 discrete actions.
@@ 4 Jointly-Optimized World-Action Model
In this section, we first detail the architecture of JOWA and the loss functions for joint optimization in section 4.1. Next, we introduce the provably efficient and parallelizable planning algorithm employed to compensate for the Q-value estimation error in section 4.2. Finally, we present the overall training pipeline in section 4.3.
@@ 4.1.2 Training of World-part Module
To unify the category of loss functions for balanced training (Vandenhende et al., 2021), we convert scalar rewards to ternary quantities $\{-1,0,1\}$ using the sign function. This allows all three predictors to be optimized as classification problems by cross-entropy loss. Given $L$-timesteps segments sampled from the offline dataset, the loss function for the world-part module is formulated as:
@@ 4.1.3 Training of Action-part Module
We use CQL for offline TD-learning during pretraining. To ensure training stability and enhance scaling performance (Farebrother et al., 2024), we employ distributional TD-error (Bellemare et al., 2017) instead of the mean-square TD-error, maintaining consistency with the $\mathcal{L}_{\text{world}}$ loss category.
The action-part module computes the return distribution for all actions given an observation and historical information. The value function $Q(o_{t},a_{t})$ is the mean of $\mathcal{Z}(o_{t},a_{t})$. Then the loss function for the action-part module is formulated as:
@@
where $\mathcal{T}^{*}$ is the distributional Bellman optimality operator defined in Equation (1), and $\mathcal{Z}^{-}$ is the target distribution computed through a target Q-values head $h_{Q^{-}}$. We set $\alpha=0.1$ in experiments.
@@ 4.2 Parallelizable planning at inference time
Given an estimated optimal Q-value $\hat{Q}^{*}$, a learned world model with dynamic predictor $\hat{P}$ and reward predictor $\hat{r}$, our objective is to find the optimal action $a^{*}_{0}$ maximizing the ground-truth optimal Q-value $Q^{*}$, starting from state $s_{0}$. To do so, we rewrite the Bellman optimality equation as:
@@
where $\pi_{Q^{*}}$ is the policy induced by the optimal Q-function. For Equation (10), the optimal policy is the greedy policy based on $Q^{*}$. The proof is provided in the Appendix A.
@@
We implement this search using beam search, an efficient decoding algorithm common in NLP. At each timestep, we retain only $K$ states with the top-$K$ total income values. The horizon $H$ and beam width $K$ are hyper-parameters, with $K=1$ or $H=0$ degenerating to a $\hat{Q}^{*}$-based greedy policy. The algorithm calls the world model $K^{2}(H-1)+K$ times but only takes $H$ times as long as the forward propagation of the world model due to parallelizability across $K$ beams.
4.3 Training pipeline
@@ 5.1 Experimental Setup
We use the Atari dataset from Agarwal et al. (2020), which contains 50M transitions from each of 5 separate training runs. Due to the prohibitive long training time on full games, we select a subset of 20 games, maintaining the difficulty distribution of full games defined by Gulcehre et al. (2020). These 20 games are introduced in Appendix D. 15 of those games are used for training, and 5 games are held out for OOD generalization experiments. Following Lee et al. (2022), we use data from 2 out of 5 training runs. To investigate performance in low-data regime, we uniformly draw 10% of transitions at random, as per Agarwal et al. (2020), resulting in 10M transitions per game.
@@
We implement our world-action model based on GPT-2 (Brown et al., 2020). We train three JOWA variants: JOWA-150M (150M parameters), JOWA-70M, and JOWA-40M. We set the number of visual tokens $K$ to 36, resulting in a total of approximately 6B tokens in our dataset. The sequence length $L$ is set to 8. We pretrain all JOWA models on A100 GPUs for 1.75M steps. For fine-tuning, we train models for 50k gradient steps with 5k transitions.
@@
During evaluation, we enable planning for JOWA, setting the planning horizon $H$ to 2 for all games and adjust the beam width $K$ based on the valid action space size of each game. See Appendix E.3 for the detailed evaluation protocol. We measure performance using human normalized scores (HNS) (Mnih et al., 2015), i.e. $(\textrm{score}-\textrm{score}_{\textrm{random}})/(\textrm{score}_{\textrm{human}}-\textrm{score}_{\textrm{random}})$. To create an aggregate comparison metric across all games, we use inter-quartile mean (IQM) of human-normalized scores, following evaluation best practices proposed in (Agarwal et al., 2021).
@@
5.2 Baseline Methods
We compare JOWA with the following baselines: (i) Multi-Task BC (MTBC). Our method can naturally be reduced to a transformer-based multi-task behavioral cloning agent. We train MTBC models at 3 scales: 34M, 65M, 120M parameters. (ii) Multi-Game DT (MGDT). Lee et al. (2022) used $\{o_{t+i},\hat{R}_{t+i},a_{t+i},r_{t+i}\}_{i=0}^{3}$ as input sequences, where $\hat{R}$ is the target return, and expert-conditioned return distribution to induce expert-level actions during evaluation. We train MGDT models at 2 scales: 40M and 200M parameters. (iii) Elastic DT (EDT). Based on MGDT, Wu et al. (2024) removed the reward in the input sequences and dynamically selected history length to address challenges of trajectory stitching. We train EDT based on the architecture configuration of MGDT-200M. (iv) Scaled-QL (SQL). Kumar et al. (2023) scaled offline Q-learning through pre-trained and shared representation network and separate Q-values head for each game. We reproduce the largest Scaled-QL with 80M parameters using ResNet-101 network. (v) FICC. Ye et al. (2022) pretrained the dynamic model with videos and then online fine-tuned the agent on each game. To obtain an offline multi-task policy, we pretrain FICC for 0.5M steps and then primarily fine-tune the $Q$-function on 15 games all at once, the same as JOWA’s second pretraining stage, for 1.25M steps. We replace all residual blocks in FICC’s official code with ResNet-50-style blocks, resulting in FICC-85M. For fair comparison, all methods use the same batch size of 512 and 1.75M gradient steps. The implementation details of all baselines are provided in the Appendix E.1.